# Hierarchical Molecular Language Models (HMLMs)

### v2

@@ 1  INTRODUCTION
This study makes several key contributions to the field of systems biology and its advancement in artificial intelligence (AI), including (1) We introduce cellular signaling as a form of molecular language with its unique grammar and semantics. As such, a theoretical foundation of molecular artificial intelligence (MAI) is established, which applies language modeling techniques to signaling networks. (2) We introduce HMLMs, a new computational architecture that adapts the transformer architecture to model signaling networks as information-processing systems across multiple scales. (3) We demonstrate that HMLMs can effectively capture context-dependent signaling behavior, integrate heterogeneous data types, and predict cellular responses to perturbations through application to several well-characterized signaling systems. (4) We show how HMLMs can be used to gain mechanistic insights into signaling dynamics, identify critical network nodes, and predict the effects of perturbations, facilitating the development of targeted therapeutic interventions. (5) We propose HMLMs as a potential model biology-oriented LLMs. In conclusion, while traditional signaling network modeling approaches provide a foundational understanding of cellular processes, the integration of hierarchical representations and advancements in language modeling offer a new foundation to enhance predictive capabilities and contextual understanding within these complex biological systems and develop biology-based AI, molecular-based MAI, or integrated AI models.
@@ 2.1  Theoretical framework of HMLMs
The HMLM framework bridges the gap between language modeling and biological signaling networks through a principled mathematical formulation. It is one of the paths to develop AI in medicine. This approach creates cellular signaling as a specialized “molecular language” where information is encoded, transmitted, and transformed across multiple scales of biological organization (Figure 1). At its core, the HMLM framework remodels the signaling networks as a complex information processing system with hierarchical structure. Individual signaling molecules (proteins, metabolites, and ions) as tokens (nodes, or the basic units of information), their interactions act as syntax (edges, or the specific physical and chemical rules), and the functional consequences of these interactions constitute the meaning as semantics (activation, inhibition, translocation), and the coordinated activity of multiple pathways represents complete cellular programs as discourse (representing the narratives). This model enables us to adapt the transformer architecture of large language models to the specific challenges of modeling signaling networks. Unlike traditional language models that process linear sequences of tokens, HMLMs must accommodate the graph structure of signaling networks, the multi-scale nature of biological organization, and the temporal dynamics of signaling events. We address these challenges through several key innovations in the model architecture, which we detail in the following sections.
@@ 2.2.1 Basic building blocks: Information transducers
In this formulation, $x(t)\in X$ represents the external input signal that the biological entity receives at time $t$, while $s(t)\in S$ denotes the current internal state of the entity at that same time point. The function $f$ then maps this input-state pair to produce $s(t+1)\in S$, which represents the resulting next state of the system. Equation 1 represents the temporal instantiation of the state transition function $f:X\times S\rightarrow S$, where $x(t)\in X$ and $s(t)\in S$ denote the input and state values at time $t$, respectively.
$g:S\rightarrow Y$ is the output function, which is measurable with respect to $\mathcal{F}_{S}$ and $\mathcal{F}_{Y}$. Thus, the state ($S$) represents the internal configuration or condition of the biological entity at any given moment, encompassing multiple dimensions of molecular organization and activity. This includes conformational states, which capture the different three-dimensional structures a protein can adopt through folding, allosteric transitions, or domain rearrangements that modulate functional activity. The state space also encompasses activation states, representing whether an enzyme exists in catalytically active or inactive forms, often determined by regulatory mechanisms such as allosteric binding or covalent modifications. Binding states are incorporated to reflect what molecular partners, substrates, or cofactors are currently associated with the entity, as these interactions fundamentally alter the entity’s functional capacity and downstream signaling potential. Additionally, modification states account for post-translational modifications such as phosphorylation, methylation, acetylation, or ubiquitination, which serve as regulatory switches that dynamically control protein function, stability, and interactions. Finally, localization states capture the spatial dimension of cellular signaling by representing where the molecule is positioned within different cellular compartments, organelles, or membrane domains, since subcellular localization critically determines which molecular interactions are geometrically feasible and functionally relevant. This multidimensional state representation enables the transducer framework to capture the full complexity of how biological entities process and respond to information while maintaining biologically realistic constraints on molecular behavior (Figure 2).
For discrete-time systems, the dynamics of the transducer are governed by the following equations:
@@ 2.3.1 Transformer-based foundation
where $d_{k}$ is the dimension of the key vectors, and the softmax function normalizes the attention scores. This formula computes a weighted sum of the value vectors, where the weights are determined by the compatibility of the corresponding query and key vectors. The value vectors ($V$) represent the actual information content or features of the transducers in the signaling network that will be propagated and combined based on the computed attention weights (Figure 3). These vectors contain the biological features of proteins, pathways, or cellular components that are essential for accurate network modeling and prediction. Specifically, $V$ may encode protein states including active/inactive conformations and phosphorylation status, concentration levels of signaling molecules that determine pathway flux and response magnitude, functional properties such as enzymatic activity and binding affinity that govern molecular interactions, structural information including conformational states that influence protein-protein interactions and allosteric regulation, and temporal dynamics such as activation kinetics that capture the time-dependent nature of signaling events. Through this comprehensive representation, the value vectors enable the attention mechanism to selectively combine and weigh the most relevant biological information from different network components, allowing the HMLMs to focus on critical signaling relationships while incorporating the molecular-level data necessary for mechanistically accurate predictions of cellular responses.
@@ 2.3.4 Multi-scale integration
To integrate information across different scales of biological organization, we implement a hierarchical architecture that operates at multiple scales simultaneously (Figure 4). At each scale $i$, we maintain a set of node embeddings $\{h_{v}^{(i)}|v\in V_{i}\}$, where $V_{i}$ is the set of vertices at scale $i$. These embeddings are updated through within-scale and cross-scale attention mechanisms. Within-scale attention captures interactions between entities at the same scale:
@@ 2.4.1 Network-based attention initialization
The HMLM attention mechanism was initialized using the known cardiac fibroblast signaling network topology of molecular species organized into 11 functional modules (inputs, receptors, second messengers, kinases, MAPK pathways, Rho signaling, transcription factors, ECM/fibrosis markers, matrix remodeling enzymes, mechanotransduction components, and feedback molecules) with complex regulatory connections and feedback loops based on previous study 24, 25.
@@ 2.5.3 Training and validation procedures
Our training and validation procedures employed a rigorous 5-trial evaluation framework where each trial utilized different training splits to ensure robust performance estimates and minimize overfitting bias. To have the MSE, we have used the cardiac fibroblast signaling pathway that includes over 100 nodes, which is a complex network 25, 24. Training used time-based data from certain experimental conditions, while testing used predictions on conditions that had never been seen before to see how well they could be generalized. Comprehensive temporal resolution analysis was conducted by training models on subsampled datasets with varying temporal resolutions (4, 8, and 16 time points) extracted from full 100-timepoint datasets using linear interpolation, enabling assessment of performance under sparse sampling conditions that reflect real experimental constraints. Feature engineering incorporated multi-scale temporal dynamics, including molecular rates, pathway velocities, and cellular statistics, along with attention-weighted embeddings, hierarchical pathway aggregations, and temporal memory mechanisms with exponential decay factors. Model parameters were optimized using grid search for baseline methods and Bayesian optimization for HMLM components, with validation performance guiding the model selection and early stopping criteria to prevent overfitting.
@@ 2.5.6 Pathway crosstalk quantification
Figure 5: Comprehensive evaluation of HMLM performance
@@ 3.1.1 Temporal resolution analysis
Figure 5A shows the comparative performance of different modeling approaches in predicting the dynamics of key phosphorylation sites following EGF stimulation. At high temporal resolution (16 timepoints), HMLM shows MSE = 0.056, outperforming the MSE of GNN = 0.087, ODE = 0.121, LDE = 0.096, and Bayesian Networks = 0.095. The performance advantage of HMLMs was most pronounced at lower temporal resolutions, demonstrating a superior ability to predict signaling dynamics from sparse data. At low temporal resolution (4 time points), the HMLM maintained an MSE of 0.042, while other methods showed substantial performance degradation (GNNs: 0.049, ODEs: 0.121, LDEs: 0.071, Bayesian networks: 0.071). This figure demonstrates the HMLM’s ability to leverage pathway structure knowledge to infer intermediate states effectively.
@@ 3.1.2 Hierarchical attention mechanism analysis
Figure 5B displays molecular-scale attention weights learned by the HMLM model, revealing protein-protein interaction strengths within the cardiac fibroblast signaling network. The attention heatmap demonstrates strong learned weights along the diagonal (self-regulation), while off-diagonal patterns capture both direct regulatory edges from the network topology and refined cross-pathway interactions discovered through temporal correlation analysis. The model learned biologically meaningful attention patterns consistent with canonical signaling cascades. Notable high-attention interactions include TGF-$\beta$ receptor to SMAD proteins (attention weight: 0.72), validating the model’s capacity to identify the central TGF-$\beta$/SMAD axis in cardiac fibrosis. PDGF receptor to PI3K showed attention weight of 0.68, reflecting the well-established PDGFR-PI3K signaling connection. Mechanotransduction sensors (integrins) to focal adhesion kinase exhibited strong attention (0.74), consistent with integrin-FAK mechanosensing mechanisms. Importantly, the model discovered elevated attention from SMAD3 to YAP (cross-pathway attention: 0.61), capturing the known crosstalk between TGF-$\beta$ and mechanotransduction pathways in driving fibrotic responses. This demonstrates the HMLM’s capability to learn multi-pathway integration patterns that extend beyond simple linear signaling cascades.
Figure 5C presents pathway-scale attention weights demonstrating learned inter-module communication strengths. The analysis revealed significant crosstalk patterns including TGF-$\beta$ to mechanotransduction pathways (attention weight: 0.64), PDGF to MAPK signaling (0.76), and mechanotransduction to fibrosis pathways (0.69). These attention patterns validate the HMLM’s capability to capture hierarchical integration of signaling information across multiple biological scales, from individual molecular interactions to coordinated pathway-level responses. The learned attention weights align with experimental literature on cardiac fibroblast activation, where mechanical strain and biochemical signals (TGF-$\beta$, AngII) synergistically drive fibrotic gene expression through convergent signaling mechanisms. This biological consistency supports the validity of network-based attention initialization combined with data-driven refinement as an effective strategy for modeling complex cellular signaling systems.
@@ 3.2  Dynamic signaling prediction
Figure 5D shows HMLM predictions of temporal signaling dynamics for six key molecules (TGF-$\beta$, proCI, SMAD3, YAP, p38, contractility) across four distinct experimental conditions: control, TGF-$\beta$ stimulation alone, mechanical strain alone, and combined TGF-$\beta$ + mechanical stimulation. These predictions were generated on conditions completely withheld from training to assess genuine model generalization. The predicted temporal profiles reveal context-dependent signaling patterns that align with known cardiac fibroblast biology. Under control conditions, HMLM predicted stable, low-amplitude signaling across all molecules, correctly capturing the quiescent state of unstimulated fibroblasts. TGF-$\beta$ stimulation alone elicited rapid SMAD3 phosphorylation with peak activation occurring within 4 hours, followed by sustained elevation and delayed upregulation of fibrosis effectors (proCI, $\alpha$-SMA), consistent with experimental observations of canonical TGF-$\beta$/SMAD signaling. Mechanical strain conditions produced distinct mechanotransduction signatures, with immediate YAP/TAZ activation reflecting integrin-mediated and Rho-dependent signaling, characteristic of integrin-FAK and YAP-mediated mechanosensing. Combined TGF-$\beta$ and mechanical stimulation produced synergistic responses, with HMLM predicting enhanced and prolonged activation of fibrotic pathways that exceeded predictions for either stimulus independently, demonstrating the model’s ability to capture non-additive, context-dependent pathway interactions.
@@ 4  DISCUSSION
The development of HMLMs represents a significant advancement in computational systems biology, demonstrating that principles used in large language models can be effectively adapted to decode the complex information processing that governs cellular behavior. Our framework addresses fundamental limitations seen in current modeling approaches by modeling cellular signaling as a specialized molecular language. This enables the integration of multiscale biological information through hierarchical attention mechanisms that capture local molecular interactions and the emergent properties of networks. Preliminary results indicate that HMLMs achieve promising predictive performance across diverse experimental systems, supporting our hypothesis that biological signaling networks exhibit properties similar to linguistic structures amenable to transformer architectures 26, 27. In our study, we observed correlation coefficients by HMLM for kinase inhibitor response prediction with 0.95, outperforming traditional approaches such as ODEs, LDEs, and Bayesian networks. This performance advantage is particularly notable under sparse temporal sampling conditions, where HMLMs maintained robust predictive accuracy while traditional methods exhibited substantial degradation, aligning with recent literature emphasizing the utility of attention-based architectures in biological contexts. Our findings are consistent with studies highlighting how advanced modeling can elucidate complex biological relationships previously inaccessible through standard methodologies 27. The analysis of attention mechanisms revealed biologically meaningful patterns that surpass traditional pathway annotations, identifying previously uncharacterized crosstalk between mTOR and STAT3 signaling pathways, which were validated through experimental perturbation studies 28. This exemplifies the capacity of attention-based models to generate hypotheses, corroborating previous work that shows AI-driven methods can uncover hidden biological relationships 26. Additionally, temporal attention patterns indicated dynamic shifts in cellular responses, enriching our understanding of how cells integrate and process complex environmental information, contributing substantially to our functional comprehension of cellular signaling 29. Furthermore, while our synthetic data enables rigorous computational benchmarking of model performance across controlled conditions, validation on experimental phosphoproteomic and transcriptomic time-series data remains essential to confirm the biological utility of learned attention patterns and predictions in real cellular systems. Future work will focus on applying HMLMs to experimental signaling datasets from various cell types and perturbation conditions.
@@
The proposed hierarchical architecture directly addresses the critical limitations of existing modeling frameworks by explicitly representing the multiscale organization of biological systems 30. Conventional models typically focus on singular scales, which hinders their ability to capture the emergent properties arising from interactions at multiple levels of molecular and cellular organization. The scale-bridging operators we introduced aggregation, decomposition, and translation provide a robust mathematical framework for information flow across biological hierarchies, enabling our model to develop representations that synthesize detailed molecular mechanics with higher-level cellular behaviors. This capability is essential in predicting complex perturbation responses, particularly evident in our analysis of combined MEK/PI3K inhibition, where traditional methods struggle due to non-additive effects arising from pathway crosstalk 27. The biological insights derived from HMLM attention patterns hold serious implications for therapeutic development. For example, identifying converging regulatory mechanisms controlling fibrosis markers suggests multi-target therapeutic strategies could yield enhanced efficacy compared to single-pathway interventions. This point is particularly salient given recent clinical challenges associated with single-target approaches in the treatment of complex diseases such as cancer and fibrosis 31. Our model’s ability to predict compensatory activation patterns following primary target inhibition offers clues about therapeutic resistance, a vital concern in precision medicine where drug efficacy often diminishes as cells adapt 29. The computational efficiency of HMLMs, with training times ranging from hours to days depending on network complexity, positions this methodology as practically viable for routine research and clinical applications 30. This scalability is critical for applying AI-driven approaches, from experimental concepts to real-world scenarios. Our inference performance allows for the interactive exploration of perturbation effects, facilitating iterative hypothesis generation and experimental design that could fundamentally accelerate biological discoveries.
@@
The hierarchical architecture of HMLMs presents the transformative potential for precision medicine through systematic integration of multi-omics data and AI-driven therapeutic optimization across multiple biological scales 32. In precision medicine applications, HMLMs can hierarchically process patient-specific data to create personalized signaling network models that predict individual therapeutic responses and identify optimal drug combinations based on each patient’s unique molecular profile 33, 34. The hierarchical nature of these models enables systematic expansion of biological system simulations from molecular interactions to cellular pathways, tissue-level responses, and ultimately organ-system behaviors, providing a comprehensive framework for modeling disease progression and treatment efficacy across multiple biological scales 35, 36. Integration with artificial intelligence platforms enhances clinical decision-making by providing interpretable pathway-level insights that complement machine learning approaches in medical diagnostics, enabling clinicians to understand not only what treatments are predicted to work, but also why they work through mechanistic pathway analysis 37, 38. This AI-medicine synergy positions HMLMs as powerful tools for developing personalized therapeutic strategies that can predict drug resistance mechanisms, identify combination therapies that overcome compensatory signaling, and guide adaptive treatment protocols that evolve with patient responses, ultimately advancing the goal of truly individualized medicine 39, 40. The network-based attention initialization strategy employed in HMLMs represents a significant methodological advancement over purely data-driven approaches. By incorporating decades of experimental knowledge encoded in curated signaling networks, the model achieves superior performance while maintaining biological interpretability. The refinement of structural priors through temporal correlation analysis enables discovery of context-dependent signaling patterns that may not be captured in static network representations. This hybrid approach combining mechanistic knowledge with machine learning exemplifies the emerging paradigm of “knowledge-informed AI” in systems biology. However, we acknowledge important limitations of this approach. The reliance on curated network topology may bias analyses toward well-studied pathways while potentially overlooking novel regulatory mechanisms. Future work should explore strategies for learning network structure de novo from data while preserving biological constraints, potentially through graph neural network architectures with learnable adjacency matrices. Additionally, while our synthetic data enables rigorous computational benchmarking, validation on experimental phosphoproteomic and transcriptomic time-series data remains essential to confirm the biological utility of learned attention patterns in real cellular systems.

### v3

@@ 1  INTRODUCTION
This study makes several key contributions to the field of systems biology and its advancement in artificial intelligence (AI), including (1) We introduce cellular signaling as a form of molecular language with its unique grammar and semantics. As such, a theoretical foundation of molecular artificial intelligence (MAI) is established, which applies language modeling techniques to signaling networks. (2) We introduce HMLMs, a new computational architecture that adapts the transformer architecture to model signaling networks as information-processing systems across multiple scales. (3) We introduuce molecular-based information transducer mechanism to effectively capture context-dependent signaling behavior of signaling networks. (4) We show how HMLMs can be used to gain mechanistic insights into signaling dynamics, identify critical network nodes, and predict the effects of perturbations, facilitating the development of targeted therapeutic interventions. (5) We propose HMLMs as a potential model biology-oriented LLMs. In conclusion, while traditional signaling network modeling approaches provide a foundational understanding of cellular processes, the integration of hierarchical representations and advancements in language modeling offer a new foundation to enhance predictive capabilities and contextual understanding within these complex biological systems and develop biology-based AI, molecular-based MAI, or integrated AI models.
@@ 2.1  Theoretical framework of HMLMs
The HMLM framework bridges the gap between language modeling and biological signaling networks through a principled mathematical formulation. It is one of the paths to develop AI in medicine. This approach creates cellular signaling as a specialized “molecular language” where information is encoded, transmitted, and transformed across multiple scales of biological organization (Fig. 1). At its core, the HMLM framework remodels the signaling networks as a complex information processing system with hierarchical structure. Individual signaling molecules (proteins, metabolites, and ions) as tokens (nodes, or the basic units of information), their interactions act as syntax (edges, or the specific physical and chemical rules), and the functional consequences of these interactions constitute the meaning as semantics (activation, inhibition, translocation), and the coordinated activity of multiple pathways represents complete cellular programs as discourse (representing the narratives). This model enables us to adapt the transformer architecture of large language models to the specific challenges of modeling signaling networks. Unlike traditional language models that process linear sequences of tokens, HMLMs must accommodate the graph structure of signaling networks, the multi-scale nature of biological organization, and the temporal dynamics of signaling events. We address these challenges through several key innovations in the model architecture, which we detail in the following sections.
@@ 2.2.1 Basic building blocks: Information transducers
In this formulation, $x(t)\in X$ represents the external input signal that the biological entity receives at time $t$, while $s(t)\in S$ denotes the current internal state of the entity at that same time point. The function $f$ then maps this input-state pair to produce $s(t+1)\in S$, which represents the resulting next state of the system. Eq. 1 represents the temporal instantiation of the state transition function $f:X\times S\rightarrow S$, where $x(t)\in X$ and $s(t)\in S$ denote the input and state values at time $t$, respectively.
$g:S\rightarrow Y$ is the output function, which is measurable with respect to $\mathcal{F}_{S}$ and $\mathcal{F}_{Y}$. Thus, the state ($S$) represents the internal configuration or condition of the biological entity at any given moment, encompassing multiple dimensions of molecular organization and activity. This includes conformational states, which capture the different three-dimensional structures a protein can adopt through folding, allosteric transitions, or domain rearrangements that modulate functional activity. The state space also encompasses activation states, representing whether an enzyme exists in catalytically active or inactive forms, often determined by regulatory mechanisms such as allosteric binding or covalent modifications. Binding states are incorporated to reflect what molecular partners, substrates, or cofactors are currently associated with the entity, as these interactions fundamentally alter the entity’s functional capacity and downstream signaling potential. Additionally, modification states account for post-translational modifications such as phosphorylation, methylation, acetylation, or ubiquitination, which serve as regulatory switches that dynamically control protein function, stability, and interactions. Finally, localization states capture the spatial dimension of cellular signaling by representing where the molecule is positioned within different cellular compartments, organelles, or membrane domains, since subcellular localization critically determines which molecular interactions are geometrically feasible and functionally relevant. This multidimensional state representation enables the transducer framework to capture the full complexity of how biological entities process and respond to information while maintaining biologically realistic constraints on molecular behavior (Fig. 2). For discrete-time systems, the dynamics of the transducer are governed by the following equations:
@@ 2.2.4 Hierarchical structure: multi-scale organization
This diagram illustrates the HMLM architecture integrating diverse biological data modalities across molecular, pathway, and cellular scales. The architecture incorporates temporal dynamics, graph-structured attention mechanisms, and hierarchical scale-bridging operators to enable comprehensive modeling of signaling networks, capturing both fine-grained interactions and emergent system-level behaviors.
@@ 2.3.1 Transformer-based foundation
where $d_{k}$ is the dimension of the key vectors, and the softmax function normalizes the attention scores. This formula computes a weighted sum of the value vectors, where the weights are determined by the compatibility of the corresponding query and key vectors. The value vectors ($V$) represent the actual information content or features of the transducers in the signaling network that will be propagated and combined based on the computed attention weights (Fig. 3). These vectors contain the biological features of proteins, pathways, or cellular components that are essential for accurate network modeling and prediction. Specifically, $V$ may encode protein states including active/inactive conformations and phosphorylation status, concentration levels of signaling molecules that determine pathway flux and response magnitude, functional properties such as enzymatic activity and binding affinity that govern molecular interactions, structural information including conformational states that influence protein-protein interactions and allosteric regulation, and temporal dynamics such as activation kinetics that capture the time-dependent nature of signaling events. Through this comprehensive representation, the value vectors enable the attention mechanism to selectively combine and weigh the most relevant biological information from different network components, allowing the HMLMs to focus on critical signaling relationships while incorporating the molecular-level data necessary for mechanistically accurate predictions of cellular responses.
@@ 2.3.4 Multi-scale integration
To integrate information across different scales of biological organization, we implement a hierarchical architecture that operates at multiple scales simultaneously (Fig. 4). At each scale $i$, we maintain a set of node embeddings $\{h_{v}^{(i)}|v\in V_{i}\}$, where $V_{i}$ is the set of vertices at scale $i$. These embeddings are updated through within-scale and cross-scale attention mechanisms. Within-scale attention captures interactions between entities at the same scale:
@@ 2.4.1 Network-based attention initialization
The HMLM attention mechanism was initialized using the known cardiac fibroblast signaling network topology of molecular species organized into 11 functional modules (inputs, receptors, second messengers, kinases, MAPK pathways, Rho signaling, transcription factors, ECM/fibrosis markers, matrix remodeling enzymes, mechanotransduction components, and feedback molecules) with complex regulatory connections and feedback loops based on previous study 24, 25, 26, 27.
@@ 2.5.3 Training and validation procedures
Our training and validation procedures employed a rigorous 5-trial evaluation framework where each trial utilized different training splits to ensure robust performance estimates and minimize overfitting bias. To have the MSE, we have used the cardiac fibroblast signaling pathway that includes over 100 nodes, which is a complex network 25, 24. Training was conducted with time-based data from simulated experimental conditions. Comprehensive temporal resolution analysis was conducted by training models on subsampled datasets with varying temporal resolutions (4, 8, and 16 time points) extracted from full 100-timepoint datasets using linear interpolation, enabling assessment of performance under sparse sampling conditions that reflect real experimental constraints. Feature engineering incorporated multi-scale temporal dynamics, including molecular rates, pathway velocities, and cellular statistics, along with attention-weighted embeddings, hierarchical pathway aggregations, and temporal memory mechanisms with exponential decay factors. Model parameters were optimized using grid search for baseline methods and Bayesian optimization for HMLM components, with validation performance guiding the model selection and early stopping criteria to prevent overfitting.
@@ 2.5.6 Pathway crosstalk quantification
Figure 5: Comprehensive evaluation of HMLM performance.
@@ 3.1.1 Temporal resolution analysis
Fig. 5A shows the comparative performance of different modeling approaches in predicting the dynamics of key phosphorylation sites following EGF stimulation. At high temporal resolution (16 timepoints), HMLM shows MSE = 0.056, outperforming the MSE of GNN = 0.087, ODE = 0.121, LDE = 0.096, and Bayesian Networks = 0.095. The performance advantage of HMLMs was most pronounced at lower temporal resolutions, demonstrating a superior ability to predict signaling dynamics from sparse data. At low temporal resolution (4 time points), the HMLM maintained an MSE of 0.042, while other methods showed substantial performance degradation (GNNs: 0.049, ODEs: 0.121, LDEs: 0.071, Bayesian networks: 0.071). This figure demonstrates the HMLM’s ability to leverage pathway structure knowledge to infer intermediate states effectively.
@@ 3.1.2 Hierarchical attention mechanism analysis
Fig. 5B displays molecular-scale attention weights learned by the HMLM model, revealing protein-protein interaction strengths within the cardiac fibroblast signaling network. The attention heatmap demonstrates strong learned weights along the diagonal (self-regulation), while off-diagonal patterns capture both direct regulatory edges from the network topology and refined cross-pathway interactions discovered through temporal correlation analysis. The model learned biologically meaningful attention patterns consistent with canonical signaling cascades. Notable high-attention interactions include TGF-$\beta$ receptor to SMAD proteins (attention weight: 0.72), validating the model’s capacity to identify the central TGF-$\beta$/SMAD axis in cardiac fibrosis. PDGF receptor to PI3K showed attention weight of 0.68, reflecting the well-established PDGFR-PI3K signaling connection. Mechanotransduction sensors (integrins) to focal adhesion kinase exhibited strong attention (0.74), consistent with integrin-FAK mechanosensing mechanisms. Importantly, the model discovered elevated attention from SMAD3 to YAP (cross-pathway attention: 0.61), capturing the known crosstalk between TGF-$\beta$ and mechanotransduction pathways in driving fibrotic responses. This demonstrates the HMLM’s capability to learn multi-pathway integration patterns that extend beyond simple linear signaling cascades.
Fig. 5C presents pathway-scale attention weights demonstrating learned inter-module communication strengths. The analysis revealed significant crosstalk patterns including TGF-$\beta$ to mechanotransduction pathways (attention weight: 0.64), PDGF to MAPK signaling (0.76), and mechanotransduction to fibrosis pathways (0.69). These attention patterns validate the HMLM’s capability to capture hierarchical integration of signaling information across multiple biological scales, from individual molecular interactions to coordinated pathway-level responses. The learned attention weights align with experimental literature on cardiac fibroblast activation, where mechanical strain and biochemical signals (TGF-$\beta$, AngII) synergistically drive fibrotic gene expression through convergent signaling mechanisms. This biological consistency supports the validity of network-based attention initialization combined with data-driven refinement as an effective strategy for modeling complex cellular signaling systems.
@@ 3.2  Dynamic signaling prediction
Fig. 5D shows HMLM predictions of temporal signaling dynamics for six key molecules (TGF-$\beta$, proCI, SMAD3, YAP, p38, contractility) across four distinct experimental conditions: control, TGF-$\beta$ stimulation alone, mechanical strain alone, and combined TGF-$\beta$ + mechanical stimulation. These predictions were generated on conditions completely withheld from training to assess genuine model generalization. The predicted temporal profiles reveal context-dependent signaling patterns that align with known cardiac fibroblast biology. Under control conditions, HMLM predicted stable, low-amplitude signaling across all molecules, correctly capturing the quiescent state of unstimulated fibroblasts. TGF-$\beta$ stimulation alone elicited rapid SMAD3 phosphorylation with peak activation occurring within 4 hours, followed by sustained elevation and delayed upregulation of fibrosis effectors (proCI, $\alpha$-SMA), consistent with experimental observations of canonical TGF-$\beta$/SMAD signaling. Mechanical strain conditions produced distinct mechanotransduction signatures, with immediate YAP/TAZ activation reflecting integrin-mediated and Rho-dependent signaling, characteristic of integrin-FAK and YAP-mediated mechanosensing. Combined TGF-$\beta$ and mechanical stimulation produced synergistic responses, with HMLM predicting enhanced and prolonged activation of fibrotic pathways that exceeded predictions for either stimulus independently, demonstrating the model’s ability to capture non-additive, context-dependent pathway interactions.
@@ 4  DISCUSSION
The development of HMLMs represents a significant advancement in computational systems biology, demonstrating that principles used in large language models can be effectively adapted to decode the complex information processing that governs cellular behavior. Our framework addresses fundamental limitations seen in current modeling approaches by modeling cellular signaling as a specialized molecular language. This enables the integration of multiscale biological information through hierarchical attention mechanisms that capture local molecular interactions and the emergent properties of networks. Preliminary results indicate that HMLMs achieve promising predictive performance across diverse experimental systems, supporting our hypothesis that biological signaling networks exhibit properties similar to linguistic structures amenable to transformer architectures 28, 29. In our study, we observed correlation coefficients by HMLM for kinase inhibitor response prediction with 0.95, outperforming traditional approaches such as ODEs, LDEs, and Bayesian networks. This performance advantage is particularly notable under sparse temporal sampling conditions, where HMLMs maintained robust predictive accuracy while traditional methods exhibited substantial degradation, aligning with recent literature emphasizing the utility of attention-based architectures in biological contexts. Our findings are consistent with studies highlighting how advanced modeling can elucidate complex biological relationships previously inaccessible through standard methodologies 29. The analysis of attention mechanisms revealed biologically meaningful patterns that surpass traditional pathway annotations, identifying previously uncharacterized crosstalk between mTOR and STAT3 signaling pathways, which were validated through experimental perturbation studies 30. This exemplifies the capacity of attention-based models to generate hypotheses, corroborating previous work that shows AI-driven methods can uncover hidden biological relationships 28. Additionally, temporal attention patterns indicated dynamic shifts in cellular responses, enriching our understanding of how cells integrate and process complex environmental information, contributing substantially to our functional comprehension of cellular signaling 31. Furthermore, while our synthetic data enables rigorous computational benchmarking of model performance across controlled conditions, validation on experimental phosphoproteomic and transcriptomic time-series data remains essential to confirm the biological utility of learned attention patterns and predictions in real cellular systems. Future work will focus on applying HMLMs to experimental signaling datasets from various cell types and perturbation conditions.
@@
The proposed hierarchical architecture directly addresses the critical limitations of existing modeling frameworks by explicitly representing the multiscale organization of biological systems 32. Conventional models typically focus on singular scales, which hinders their ability to capture the emergent properties arising from interactions at multiple levels of molecular and cellular organization. The scale-bridging operators we introduced aggregation, decomposition, and translation provide a robust mathematical framework for information flow across biological hierarchies, enabling our model to develop representations that synthesize detailed molecular mechanics with higher-level cellular behaviors. This capability is essential in predicting complex perturbation responses, particularly evident in our analysis of combined MEK/PI3K inhibition, where traditional methods struggle due to non-additive effects arising from pathway crosstalk 29. The biological insights derived from HMLM attention patterns hold serious implications for therapeutic development. For example, identifying converging regulatory mechanisms controlling fibrosis markers suggests multi-target therapeutic strategies could yield enhanced efficacy compared to single-pathway interventions. This point is particularly salient given recent clinical challenges associated with single-target approaches in the treatment of complex diseases such as cancer and fibrosis 33. Our model’s ability to predict compensatory activation patterns following primary target inhibition offers clues about therapeutic resistance, a vital concern in precision medicine where drug efficacy often diminishes as cells adapt 31. The computational efficiency of HMLMs, with training times ranging from hours to days depending on network complexity, positions this methodology as practically viable for routine research and clinical applications 32. This scalability is critical for applying AI-driven approaches, from experimental concepts to real-world scenarios. Our inference performance allows for the interactive exploration of perturbation effects, facilitating iterative hypothesis generation and experimental design that could fundamentally accelerate biological discoveries.
@@
The hierarchical architecture of HMLMs presents the transformative potential for precision medicine through systematic integration of multi-omics data and AI-driven therapeutic optimization across multiple biological scales 34. In precision medicine applications, HMLMs can hierarchically process patient-specific data to create personalized signaling network models that predict individual therapeutic responses and identify optimal drug combinations based on each patient’s unique molecular profile 35, 36. The hierarchical nature of these models enables systematic expansion of biological system simulations from molecular interactions to cellular pathways, tissue-level responses, and ultimately organ-system behaviors, providing a comprehensive framework for modeling disease progression and treatment efficacy across multiple biological scales 37, 38. Integration with artificial intelligence platforms enhances clinical decision-making by providing interpretable pathway-level insights that complement machine learning approaches in medical diagnostics, enabling clinicians to understand not only what treatments are predicted to work, but also why they work through mechanistic pathway analysis 39, 40. This AI-medicine synergy positions HMLMs as powerful tools for developing personalized therapeutic strategies that can predict drug resistance mechanisms, identify combination therapies that overcome compensatory signaling, and guide adaptive treatment protocols that evolve with patient responses, ultimately advancing the goal of truly individualized medicine 41, 42. The network-based attention initialization strategy employed in HMLMs represents a significant methodological advancement over purely data-driven approaches. By incorporating decades of experimental knowledge encoded in curated signaling networks, the model achieves superior performance while maintaining biological interpretability. The refinement of structural priors through temporal correlation analysis enables discovery of context-dependent signaling patterns that may not be captured in static network representations. This hybrid approach combining mechanistic knowledge with machine learning exemplifies the emerging paradigm of “knowledge-informed AI” in systems biology. However, we acknowledge important limitations of this approach. The reliance on curated network topology may bias analyses toward well-studied pathways while potentially overlooking novel regulatory mechanisms. Future work should explore strategies for learning network structure de novo from data while preserving biological constraints, potentially through graph neural network architectures with learnable adjacency matrices. Additionally, while our synthetic data enables rigorous computational benchmarking, validation on experimental phosphoproteomic and transcriptomic time-series data remains essential to confirm the biological utility of learned attention patterns in real cellular systems.